3.2g of pyrolusite was treated with 50ml of 0.5 M oxalic acid and some sulphuric acid. The oxalic acid left undecomposed was raised to 250ml in a flask. 25 ml of this solution when treated with 0.02M kmno4 required 32ml of the solution. find the % of mno2 in the sample and also the % of available oxygen

Grade:11

1 Answers

Ravi Kumar
4 years ago
Mill eq of kmno4=0.02*32*5=3.2  So eq of italic acid=3.2in 25 ml so in 250ml=32.    WT of mno2= 50-32/1000*87/2=.783(converted Mill equal to equal and 87 is mass of mno2 and nf is 2)    percentage of mno2=.783/3.2*100=24.46